# A 4 - 8 GHz Kinetic Inductance Travelling-Wave Parametric Amplifier Using Four-Wave Mixing with Near Quantum-Limit Noise Performance

### v3

@@ Abstract
Kinetic inductance traveling-wave parametric amplifiers (KI-TWPA) have a wide instantaneous bandwidth with near quantum-limited sensitivity and a relatively high dynamic range. Because of this, they are suitable readout devices for cryogenic detectors and superconducting qubits and have a variety of applications in quantum sensing. This work discusses the design, fabrication, and performance of a KI-TWPA based on four-wave mixing in a NbTiN microstrip transmission line. This device amplifies a signal band from 4 to 8 GHz without contamination from image tones, which are produced in a separate higher frequency band. The 4 - 8 GHz band is commonly used to read out cryogenic detectors, such as microwave kinetic inductance detectors (MKIDs) and Josephson junction-based qubits. We report a measured maximum gain of over 20 dB using four-wave mixing with a 1-dB gain compression point of -58 dBm at 15 dB of gain over that band. The bandwidth and peak gain are tunable by adjusting the pump-tone frequency and power. Using a Y-factor method, we measure an amplifier-added noise of $0.5\leq N_{added}\leq 1.5$ photons from 4.5 - 8 GHz.
@@ I Introduction
Using a quantum noise-limited parametric amplifier as a first-stage amplifier for readout can improve the sensitivity of multiple cryogenic detector technologies, such as the Microwave Kinetic Inductance Detector (MKID) Zobrist et al. (2019), , and microwave SQUID multiplexer ($\mathrm{\mu}$MUX) readout of Mettalic Magnetic Calorimeters Kempf et al. (2017) and Transition-Edge Sensors Malnou et al. (2023). Traveling wave parametric amplifiers are also an appealing choice for fast and high-fidelity readout Peng et al. (2022); Barzanjeh, DiVincenzo, and Terhal (2014); Didier et al. (2015) of cryogenic qubits. A commonly used frequency band to read out superconducting qubits and detectors is 4 - 8 GHz due to its availability, accessibility, and maturity of readout electronic systems and components at a relatively low cost. Applications of superconducting parametric amplifiers also extend to fundamental physics research for laboratory-based experiments such as dark matter searches, where a quantum-noise-limited gain is of interest Ramanathan et al. (2023).
The resonant Josephson junction amplifiers (JPAs) are the most commonly used superconducting parametric amplifiers Aumentado (2020). Even though these devices have shown quantum-limited noise performance, they have low fractional bandwidths and very low dynamic rangeEsposito et al. (2021). Recently, by implementing a traveling-wave periodic structureAumentado (2020); Mutus et al. (2013), the bandwidths of Josephson traveling-wave parametric amplifiers (JTWPAs) have been increased to a few gigahertz Macklin et al. (2015). However, the low dynamic range of these devices still remains an issue, especially for the readout of resonator arrays Mutus et al. (2013). In addition, the complexity of circuitry of the nonlinear lumped element transmission line and junction fabrication also puts them at a disadvantage compared to their counterparts, the kinetic inductance-based parametric amplifiersShu et al. (2021); Malnou et al. (2021); Eom et al. (2012); Ranzani et al. (2018); Vissers et al. (2016); Chaudhuri et al. (2017); Shan, Sekimoto, and Noguchi (2016); Boc (2014).
@@
This paper presents a degenerate four-wave mixing KI-TWPA, where the pump tone is between the signal and idler frequencies, $2f_{pump}=f_{signal}+f_{idler}$. The amplifier tends to produce a gain in disjointed frequency ranges that we will refer to as the signal and idler (or image) bands, which makes it possible to separate the idler using ancillary circuitry and use the full signal band without contamination from the image (or idler) frequencies. The devices described here are made of NbTiN and patterned in an inverted microstrip geometryShu et al. (2021). The transmission line is dispersion engineered by adding sinusoidally modulated capacitive stubs with a median length of 9.2 $\mathrm{\mu m}$ and modulation of 2.08 $\mathrm{\mu m}$ to create a bandgap centered at 12.5 GHz.
@@ II Device Design and Fabrication
The design of the device is similar to what was reported in Shu et al. Shu et al. (2021) with the exception of Nb as the ground plane. To effectively increase the nonlinearity in the kinetic inductance of the transmission line as a function of the pump current ($I_{p}$), the transmission line width was designed to be 320 nm as shown in Fig. 1.a. The kinetic inductance per unit length as a function of pump current is given by the following relationEom et al. (2012):
@@
The increase in the impedance of the transmission line, $Z\approx\sqrt{\mathcal{L}_{k}/\mathcal{C}}$, due to its narrow width and large kinetic inductance was compensated for by adding capacitive stubs to increase the capacitance per unit length,$\mathcal{C}$, of the transmission line. The length of the stubs was set to create the dispersion needed for the four-wave mixing process and further modulated along the length of the transmission line to create a photonic bandgap Eom et al. (2012) around the pump frequency, as shown in Fig 1.b.
@@
$\Delta\beta=\kappa_{s}+\kappa_{i}-2\kappa_{p}=-\kappa_{p}\;\frac{I_{p}^{2}}{4I_{*}^{2}}\,,$ (2)
@@
Figure 1: a) A SEM micrograph of the transmission line with capacitive stubs. The width of the stubs is 340 nm, the spacing between them is 2.2 microns, and the average length of the stubs is 10.8 microns. b) Plot of the calculated propagation constant with a linear part subtracted as a function of frequency with a discontinuity at around 12.5 GHz, which results from the modulation of the length of the capacitive stubs with a periodicity of 122.7 microns. The inset shows the structure of the simulated dispersion around the bandgap and the position of the pump tone relative to the bandgap. c) Plot of calculated phase matching criterion, $\Delta\beta$, as a function of frequency. Shown in red is the self-phase modulation of the pump,$\kappa_{p}\frac{I_{p}^{2}}{4I_{*}^{2}}$, and in purple is the difference between the pump wave number, $2\kappa_{p}$, relative to the signal and idler wave numbers, $\kappa_{s}$ and $\kappa_{i}$. The optimal parametric gain (maximum gain) is achieved near the outer intersecting points.
The devices were fabricated by depositing a 35-nm thick NbTiN layer on a 6-inch high resistivity ($\rho\geq 10,000\;\mathrm{\Omega.cm}$) silicon wafer via reactive sputtering from a NbTi target in a nitrogen atmosphere at ambient temperatures. The NbTiN film was then patterned using a stepper photo-lithography method, and then etched in a reactive ion etcher. Utilizing a Plasma-Enhanced Chemical Vapor Deposition (PECVD) process, a 60-nm thick amorphous silicon layer was deposited on top of the NbTiN wire. Finally, a 350-nm thick Nb ground-plane layer was sputtered on top of the amorphous silicon layer and then patterned and etched.
Figure 2: a) Shows the measured gain of the device at 1 K with pump frequency at 10.6 GHz. The frequencies over which the device produces gain are split into three bands. The gap at around 12 GHz is due to the “engineered” band gap produced by the periodic modulation of the transmission line structureShu et al. (2021); Klimovich (2022). The gap around 8 GHz corresponds to frequencies where the idler tone ends up in the band gap, so the 4WM process is inoperative. The dashed line shows the simulated gained curve with the same pump condition as the measured gain of the device. b) plot of measured gain of the device at 1 K with different pump frequencies. The bandwidth of the “signal” band can be adjusted significantly by varying the pump frequency. Measurements shown in this sub-figure are from another device on the same wafer with an identical design.
@@ III.1 Behavior
The expected gain of the device was calculated using the low power dispersion extracted from the measured unpumped transmission. The measurement was compared to a network model of the KI-TWPA to determine the film surface inductance. In our method, described in ref. Shu et al. (2021), the modulated transmission line is considered to be a uniform medium with an effective dielectric constant, which we derive from the adjusted network model. That effective dielectric constant is used in the coupled-mode equations to predict the gain. For the simulated gain curve shown in Fig. 2.a, the higher-frequency products at $2\omega_{p}+\omega_{s}$, $3\omega_{p}$, and $4\omega_{p}-\omega_{s}$ were included in the calculation in addition to the signal, pump and idler frequenciesShu et al. (2021). To achieve a similar gain level to the measured value, we adjusted the value of the pump current used in the calculation. Comparing the simulation plot shown in Fig. 2.a indicates a good agreement with measurement results.
@@
The frequency band over which the device produces gain can be tuned significantly by adjusting the frequency of the pump tone, as shown in Fig. 2.b. As the dispersion introduced by the periodic modulation changes rapidly below the bandgap (Fig. 1.b), small pump frequency changes result in somewhat larger changes in the signal and idler frequencies that satisfy the phase matching condition, eqn. 2.
@@
To measure gain compression, a test tone in the middle of the device’s signal band was injected into the KI-TWPA, and the system’s output power was measured using a spectrum analyzer. With the pump frequency set to 10.6 GHz and pump tone power of -23 dBm, resulting in 15 dB of gain, the output power of the amplified test tone at the spectrum analyzer was measured as a function of the test tone’s power at the device. As shown in Fig. 3.a, the gain compresses by 1 dB when the input signal is $\simeq$ -58 dBm, which corresponds to an output power of -43 dBm. As the power of the output signal and idler tones is provided by the pump, compression occurs when the pump is level significantly depleted. For our device, that occurs with an output signal power of about -20 dB below the pump power.
@@
Figure 3: a) Plot of the measured output power at the spectrum analyzer of a test tone as a function of the test tone’s input power at 20 mK with a signal gain of 15 dB. The gain compresses by 1 dB from the linear response when the input signal power reaches -58 dBm. b) Plot of the measured gain for a fixed pump tone frequency as a function of pump tone power at 20 mK. Both of these measurements were taken using the setup shown in Fig. 5. The higher ripple level in this plot compared to Fig. 2.a is due to the addition of more components in the measurement setup, such as the diplexers and more cabling, which can cause impedance mismatch.
@@
Figure 4: a) Schematic of the noise measurement setup with hot and cold broadband noise sources. Two 50 $\mathrm{\Omega}$ cryogenic loads were used as noise sources. The hot source was heat sunk to the 3 K stage of the dilution fridge with a base temperature of 3.18 K, and the cold noise source was placed on the fridge’s mixing chamber (MXC) with a base temperature of $\simeq 20$ mK. All other components in this figure were placed in the MXC. Using a relay switch S, the input of the KI-TWPA was switched between the hot and cold sources. The two types of diplexers have crossovers in the lower and upper gaps in the gain curve shown in Fig. 2.a and are used to inject and separate out the pump and idler tones. c) circuit diagram of the cascaded noise model for the KI-TWPA measurement setup.
@@ III.2 Noise
The amplifier chain noise was measured using a Y-factor method with two 50 $\mathrm{\Omega}$ cryogenic loads as noise sources, one at 3.18 K and the other at the Mixing Chamber (MXC) stage ($\sim$ 20 mK) of the dilution refrigerator. Using a cryogenic relay switch, we alternate between the noise sources as shown in Fig. 4.a. Due to a lack of output-to-input isolation, reflections must be avoided at either side of the TWPA at frequencies where the amplifier has gain. If the sum of the dB return losses on the input and output sides does not exceed the gain, an oscillation will develop. The circuit shown in Fig. 4 provides an impedance match at both the input and output of the KI-TWPA over the signal and idler frequency ranges using a combination of diplexer (DPX) circuits. The input side diplexers allow for the injection of the pump tone through a bandpass filter (not shown) that removes synthesizer phase noise. The input side DPX I provides a cold termination to the input of the KI-TWPA over the idler band ($>$13 GHz). The noise from that termination determines the added noise of the KI-TWPA, so it is kept at a temperature $T\sim 20\,$mK $\ll\hbar\omega_{i}/k_{B}$ for the system to be quantum limited. In this configuration, the KI-TWPA only receives the thermal noise emitted from the calibration sources in the signal band.
@@
where $L_{1}$ and $L_{2}$ are the losses in the components before and after the KI-TWPA, respectively. $N_{qm}=0.5$ is the quantum noise, $N_{HEMT}$ is the added noise of the HEMT amplifier and $G_{PA}$ is the measured gain of the KI-TWPA. The noise model used to arrive at equation (6) is explained in Appendix B. We measure a total system noise of $1\leq N_{sys}\leq 2$ in the 4.6 - 8.1 GHz range as shown in Fig. 4.c. Using equation (6), we estimate the added noise of the KI-TWPA to be $0.5\leq N_{sys}\leq 1.5$ as shown in Fig. 4. The cryogenic HEMT amplifier’s noise was measured using the Y-factor with the pump off. We measured an average noise temperature of 5 K or $\sim$ 13 quanta in the 4 - 8 GHZ band consistent with data provided in the amplifier’s datasheet. It was observed that the measured added noise in the KI-TWPA deviates from half a quanta in some places (Fig. 4.c). These discrepancies could be attributed to some nonidealities in the test setup. For instance, both noise sources are connected to the input of the KI-TWPA through three cables and two switches. Estimating the insertion loss in all these components is quite challenging, and any unaccounted losses can lead to uncertainty in the measurement and overestimation of the measured noise. Impedance mismatch between the noise terminations and the KI-TWPA is also another source of uncertainty in our noise measurements. The ripples in the measured system noise and the TWPA-added noise are indications of these imperfections.

### v4

@@ Abstract
Kinetic inductance traveling-wave parametric amplifiers (KI-TWPA) have a wide instantaneous bandwidth with near quantum-limited noise performance and a relatively high dynamic range. Because of this, they are suitable readout devices for cryogenic detectors and superconducting qubits and have a variety of applications in quantum sensing. This work discusses the design, fabrication, and performance of a KI-TWPA based on four-wave mixing in a NbTiN microstrip transmission line. This device amplifies a signal band from 4 to 8 GHz without contamination from image tones, which are produced in a separate higher frequency band. The 4 - 8 GHz band is commonly used to read out cryogenic detectors, such as microwave kinetic inductance detectors (MKIDs) and Josephson junction-based qubits. We report a measured maximum gain of over 20 dB using four-wave mixing with a 1-dB gain compression point of -58 dBm at 15 dB of gain over that band. The bandwidth and peak gain are tunable by adjusting the pump-tone frequency and power. Using a Y-factor method, we measure an amplifier-added noise of $0.5\leq N_{added}\leq 1.5$ photons from 4.5 - 8 GHz.111© 2024. All rights reserved
@@ I Introduction
Using a quantum noise-limited parametric amplifier as a first-stage amplifier for readout can improve the sensitivity of multiple cryogenic detector technologies, such as the Microwave Kinetic Inductance Detector (MKID) Zobrist et al. (2019), and microwave SQUID multiplexer ($\mathrm{\mu}$MUX) readout of Metallic Magnetic Calorimeters Kempf et al. (2017) and Transition-Edge Sensors Malnou et al. (2023). Traveling wave parametric amplifiers are also an appealing choice for fast and high-fidelity readout Peng et al. (2022); Barzanjeh, DiVincenzo, and Terhal (2014); Didier et al. (2015) of cryogenic qubits. A commonly used frequency band to read out superconducting qubits and detectors is 4 - 8 GHz due to its availability, accessibility, and maturity of readout electronic systems and components at a relatively low cost. Applications of superconducting parametric amplifiers also extend to fundamental physics research for laboratory-based experiments such as dark matter searches, where a quantum-noise-limited gain is of interest Ramanathan et al. (2023).
The resonant Josephson junction amplifiers (JPAs) are the most commonly used superconducting parametric amplifiers Aumentado (2020). Even though these devices have shown quantum-limited noise performance, they have low fractional bandwidths and very low dynamic rangeEsposito (2021). Recently, by implementing a traveling-wave periodic structureAumentado (2020); Mutus et al. (2013), the bandwidths of Josephson traveling-wave parametric amplifiers (JTWPAs) have been increased to a few gigahertz Macklin et al. (2015). However, the low dynamic range of these devices still remains an issue, especially for the readout of resonator arrays Mutus et al. (2013). In addition, the complexity of circuitry of the nonlinear lumped element transmission line and junction fabrication also puts them at a disadvantage compared to their counterparts, the kinetic inductance-based parametric amplifiersShu et al. (2021); Malnou et al. (2021); Eom et al. (2012); Ranzani et al. (2018); Vissers et al. (2016); Chaudhuri et al. (2017); Shan, Sekimoto, and Noguchi (2016); Bockstiegel et al. (2014).
@@
This paper presents a degenerate four-wave mixing KI-TWPA, where the pump tone is between the signal and idler frequencies, $2f_{pump}=f_{signal}+f_{idler}$. The amplifier tends to produce gain in disjointed frequency ranges that we will refer to as the signal and idler (or image) bands, which makes it possible to separate the idler using ancillary circuitry and use the full signal band without contamination from the image (or idler) frequencies. The devices described here are made of NbTiN as the high kinetic inductance material and patterned in an inverted microstrip geometryShu et al. (2021).
@@ II Device Design and Fabrication
The design of the device is similar to what was reported in Shu et al. Shu et al. (2021) with the exception of Nb as the ground plane. To achieve a high nonlinearity in the kinetic inductance of the transmission line as a function of the pump current ($I_{p}$), the transmission line width was designed to be 340 nm as shown in Fig. 1.a. The kinetic inductance per unit length as a function of pump current is given by the following relationEom et al. (2012):
@@
The increase in the impedance of the transmission line, $Z\approx\sqrt{\mathcal{L}_{k}/\mathcal{C}}$, due to its narrow width and large kinetic inductance was compensated for by adding capacitive stubs to increase the capacitance per unit length,$\mathcal{C}$, of the transmission line. The total length of the device is 100 mm, and the calculated propagation velocity and the characteristic impedance are 0.010c and 50 $\mathrm{\Omega}$, respectively. Using these values, we estimate an inductance per unit length of $\mathcal{L}=16.64\;\mathrm{\mu H/m}$ and a capacitance per unit length of $\mathcal{C}=6.45\;\mathrm{nF/m}$. The length of the stubs was set to create the dispersion needed for the four-wave mixing process and further modulated along the length of the transmission line to create a photonic bandgap Eom et al. (2012) around the pump frequency, as shown in Fig 1.b. The average length of the capacitive stubs is 10.8 $\mathrm{\mu m}$ with amplitude modulation of 2.08 $\mathrm{\mu m}$ and periodic modulation of 122 $\mathrm{\mu m}$ to create a bandgap centered at 12.5 GHz.
@@
$\Delta\beta\equiv\kappa_{s}+\kappa_{i}-2\kappa_{p}=-\kappa_{p}\;\frac{I_{p}^{2}}{4I_{*}^{2}}\,,$ (2)
@@
Figure 1: a) A SEM micrograph of the transmission line with capacitive stubs. The width of the stubs is 340 nm, the spacing between them is 2.2 microns, and the average length of the stubs is 10.8 microns. b) Plot of k, the calculated propagation constant, with a linear part subtracted. The discontinuity at around 12.5 GHz results from the modulation of the length of the capacitive stubs with a periodicity of 122.7 microns. The red arrow shows the position of the pump tone relative to the bandgap. c) Plot of calculated phase matching criterion, $\Delta\beta$, as a function of frequency. The red line is the self-phase modulation of the pump, $\kappa_{p}I_{p}^{2}/(4I_{*}^{2})$. The purple line is $\kappa_{s}+\kappa_{i}-2\kappa_{p}$. The optimal parametric gain (maximum gain) is achieved near the outer intersecting points.
The devices were fabricated by depositing a $35\pm 5$ nm thick NbTiN layer on a 6-inch high resistivity ($\rho\geq 10,000\;\mathrm{\Omega.cm}$) silicon wafer via reactive sputtering from a NbTi target in a nitrogen atmosphere at ambient temperatures. The NbTiN film was then patterned using a stepper photo-lithography method and etched in a reactive ion etcher. Utilizing a Plasma-Enhanced Chemical Vapor Deposition (PECVD) process, a 100-nm thick amorphous silicon layer was deposited on top of the NbTiN wire. Finally, a 350-nm thick Nb ground-plane layer was sputtered on top of the amorphous silicon layer and then patterned and etched.
Figure 2: a) Measured gain of the device at 1K using a simplified cold measurement setup consisting of only a 20 dB attenuator at 1K before the KI-TWPA and input and output transmission lines directly to room temperature. The 10.6 GHz pump tone is injected through a directional coupler at room temperature. The frequencies over which the device produces gain are split into three bands. The gap at around 12.5 GHz is due to the “engineered” band gap produced by the periodic modulation of the transmission line structureShu et al. (2021); Klimovich (2022). The dashed line shows the simulated gain curve and the inset shows the gain ripples in a narrow frequency range. b) Plot of measured gain of the device at 1 K with different pump frequencies. The bandwidth of the “signal” band can be adjusted significantly by varying the pump frequency. Measurements shown in this sub-figure are from another device on the same wafer with an identical design.
@@ III.1 Behavior
The expected gain of the device was calculated using the low power dispersion extracted from the measured unpumped transmission. The measurement was compared to a network model of the KI-TWPA to determine the film surface inductance. In our method, described in ref. Shu et al. (2021), the modulated transmission line is considered to be a uniform medium with an effective dielectric constant, which we derive from the adjusted network model. That effective dielectric constant is used in the coupled-mode equations to predict the gain. For the simulated gain curve shown in Fig. 2.a, the higher-frequency products at $2\omega_{p}+\omega_{s}$, $3\omega_{p}$, and $4\omega_{p}-\omega_{s}$ were included in the calculation in addition to the signal, pump, and idler frequenciesShu et al. (2021). To match the measurement, we set the pump current in the simulation to $I_{P}=$380 $\mathrm{\mu A}$. The $I_{\ast}$ value used was 3.2 mA, which was determined by measuring the shift in the stopband as a function of DC current similar to the measurements in Shu et al.  Shu et al. (2021), which results in a ratio of the pump current to the characteristic current of $I_{p}/I_{*}=0.119$. The pump power for the measured gain curve in Fig. 2.a is -20.5 $\pm$ 1.5 dBm, where the error estimate comes from uncertainty in the cryogenic losses in the input line. The corresponding root-mean-square current is $I_{p}=421\pm 70\;\mathrm{\mu A}$, close to the value used for the simulation. Some discrepancy between those values may be expected because our model neglects the weakly standing wave nature of the pump that arises from the periodic modulation of the structure. The close agreement between the overall shape and magnitude of the measured and simulated gain curves shown in the figure nevertheless demonstrates that the model captures the main processes occurring in the device and has useful predictive properties.
@@
The frequency band over which the device produces gain can be tuned significantly by adjusting the frequency of the pump tone, as shown in Fig. 2.b. As the dispersion introduced by the periodic modulation changes rapidly below the bandgap (Fig. 1.b), small pump frequency changes result in somewhat larger changes in the signal and idler frequencies that satisfy the phase matching condition, equation (2).
The measured gain has a rapid frequency variation, shown in Fig. 2.a . The spacing between the gain ripples is 15 MHz, corresponding to a free space length of $\sim$ 20 m, which is twice the electrical length of the KI-TWPA. Therefore, we attribute the ripples to impedance mismatches between the KI-TWPA chip and the feedline circuitry on both sides of the device. Fig. 3.b shows the measured gain as a function of pump power with the KI-TWPA inserted in the noise measurement setup of Fig. 5. Compared to Fig. 2, the ripple amplitude is higher due to the return losses of the additional components. The increase in the ripple amplitude with pump power can be understood by considering a medium with one-way amplitude gain $g$ between two discontinuities represented by reflection and transmission coefficients $r$ and $t$. The forward transmission of the structure is
$S_{21}=\frac{t^{2}ge^{-ikL}}{1-r^{2}ge^{-ikL}},$ (3)
where $k$ is the propagation constant in the medium and $L$ is the length. The ripple level increases with $g$ and diverges at $g=r^{-2}$.
@@
To measure gain compression, a test tone in the middle of the device’s signal band was injected into the KI-TWPA, and the system’s output power was measured using a spectrum analyzer. With the pump frequency set to 10.6 GHz and pump tone power of -23 dBm, resulting in 15 dB of gain, the output power of the amplified test tone at the spectrum analyzer was measured as a function of the test tone’s power at the device. As shown in Fig. 3.a, the gain compresses by 1 dB when the input signal is $\simeq$ -58 dBm, corresponding to an output power of -43 dBm. As the power of the output signal and idler tones is provided by the pump, compression occurs when the pump level is significantly depleted. For our device, that occurs with an output signal power of about 20 dB below the pump power.
@@
Figure 3: a) Plot of the measured output power at the device of a test tone as a function of the test tone’s input power at 20 mK with a signal gain of 15 dB. The gain compresses by 1 dB from the linear response when the input signal power reaches -58 dBm. b) Plot of the measured gain for a fixed pump tone frequency as a function of pump tone power at 20 mK. Both of these measurements were taken using the setup shown in Fig. 5. The higher ripple level in this plot compared to Fig. 2.a is due to the addition of more components in the measurement setup, such as the diplexers and more cabling, which can cause impedance mismatch.
@@
Figure 4: a) Schematic of the noise measurement setup with hot and cold broadband noise sources. Two 50 $\mathrm{\Omega}$ cryogenic loads were used as noise sources. The hot source was heat sunk to the 3 K stage of the dilution fridge with a base temperature of 3.18 K, and the cold noise source was placed on the fridge’s mixing chamber (MXC) with a base temperature of $\simeq 20$ mK. All other components in this figure were placed in the MXC. Using a relay switch S, the input of the KI-TWPA was switched between the hot and cold sources. The two types of diplexers have crossovers in the lower and upper gaps in the gain curve shown in Fig. 2.a and are used to inject and separate out the pump and idler tones. b) circuit diagram of the cascaded noise model for the KI-TWPA measurement setup. c) Plot of the system noise and KI-TWPA added noise in units of quanta in the 4.6 - 8 GHz range. The standard quantum limit (SQL) is shown with a dashed line.
@@ III.2 Noise
The amplifier chain noise was measured using a Y-factor method with two 50 $\mathrm{\Omega}$ cryogenic loads as noise sources, one at 3.18 K and the other at the Mixing Chamber (MXC) stage ($\sim$ 20 mK) of the dilution refrigerator. Using a cryogenic relay switch, we alternate between the noise sources as shown in Fig. 4.a. Due to a lack of output-to-input isolation, reflections must be avoided at either side of the TWPA at frequencies where the amplifier has gain. If the sum of the dB return losses on the input and output sides does not exceed the gain, an oscillation will develop. The circuit shown in Fig. 4 provides an impedance match at both the input and output of the KI-TWPA over the signal and idler frequency ranges using a combination of diplexer (DPX) circuits. The input side diplexers allow for the injection of the pump tone through a bandpass filter (not shown) that removes room temperature thermal noise and synthesizer phase noise in signal and idler bands that the diplexers may not sufficiently attenuate. The input side DPX 1 provides a cold termination to the input of the KI-TWPA over the idler band ($>$13 GHz). The noise from that termination determines the added noise of the KI-TWPA, so it is kept at a temperature $T\sim 20\,$mK $\ll\hbar\omega_{i}/k_{B}$ for the system to be quantum limited. In this configuration, the KI-TWPA only receives the thermal noise emitted from the calibration sources in the signal band.
@@
where $L_{1}$ and $L_{2}$ are the losses in the components before and after the KI-TWPA, respectively. $N_{qm}=0.5$ is the quantum noise, $N_{HEMT}$ is the added noise of the HEMT amplifier and $G_{PA}$ is the measured gain of the KI-TWPA. The noise model used to arrive at equation (6) is explained in Appendix B. We measure a total system noise of $1\leq N_{sys}\leq 3$ in the 4.6 - 8 GHz range as shown in Fig. 4.c. Using equation (6), we estimate the added noise of the KI-TWPA to be $0.5\leq N_{A}\leq 1.5$ as shown in Fig. 4. The cryogenic HEMT amplifier’s noise was measured using the Y-factor with the pump off. We measured an average noise temperature of 5 K or $\sim$ 13 quanta in the 4 - 8 GHz band consistent with data provided in the amplifier’s datasheet. It was observed that the measured added noise in the KI-TWPA deviates from half a quanta in some places (Fig. 4.c). These discrepancies could be attributed to some nonidealities in the test setup. For instance, both noise sources are connected to the input of the KI-TWPA through three cables and two switches. Estimating the insertion loss in all these components is quite challenging, and any unaccounted losses can lead to uncertainty in the measurement and overestimation of the measured noise. Impedance mismatch between the noise terminations and the KI-TWPA is also another source of uncertainty in our noise measurements. The ripples in the measured system noise and the TWPA-added noise are indications of these imperfections.